the value of cos12.cos24.cos36.cos48.cos72.cos84=?a)1/64. b)1/32c)1/16. d)1/128

Grade:11

1 Answers

harsha
6 years ago
cosa*cos(60+a*)cos(60-a)=1/4cos3a then. cos12*cos48*cos72=cos12*cos(60-12)*cos(60+12)=1/4*cos(36)=(root5)+1/16eqn1and cos24*cos36*cos84=1/4(cos72)=(root5)-1/16 eqn2eqn1*eqn2=1/64